OpenStudy (anonymous):

Indicate the equation of the line that is the perpendicular bisector of the segment with endpoints (4, 1) and (2, -5).

OpenStudy (anonymous):

heelllpppp pleaseeee

OpenStudy (superhelp101):

A perpendicular bisector of a side of a triangle is a line segment that is both perpendicular to a side of a triangle and passes through its midpoint.

OpenStudy (anonymous):

how would it be written out though ?

OpenStudy (anonymous):

Do you know how to write the equation of a line pass through 2 given points? Do you know how to find the midpoint of a segment? Do you know how to write the equation of a line perpendicular with a given line? If you say yes for all, you can solve this problem :)

OpenStudy (superhelp101):

you have to find the mid- point mid-point of (x1,y1) and (x2,y2) x= (x1+x2)/2 y= (y1+y2)/2

OpenStudy (aum):

Find the slope of the line: Slope = \(\Large \frac{y_2 - y_1}{x_2-x_1}\) Slope 'm' of the perpendicular line will be the negative reciprocal of the above. Find the midpoint of the line: \(\Large (\frac{x_1+x_2}{2}, \frac{y_1+y_2}{2})\) So you know the slope of the perpendicular line and you know it passes through the midpoint. The equation of the line is: \(\Large y-y_1 = m(x-x_1)\)

OpenStudy (anonymous):

i just need a quick answer i am running out of time. thaannkksss (:

OpenStudy (anonymous):

i know the formulas i just need to know the answer

OpenStudy (aum):

Sorry. Can't give out answers. I can help you arrive at the answer but you will have to do the problem yourself.